Actions
20160126 » History » Revision 16
« Previous |
Revision 16/22
(diff)
| Next »
Boyd, Steven, 27 January 2016 07:54
Commissioning and Ops Scheduling meeting : 26/1/2016¶
To be held in CR10 in the ATLAS Building @ 15:30 pm
Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402
Agenda:¶
- Review of tasks and activities identified at last meeting
- Realignment work
- Action Item (Colin to prod John C) : need to get John's data with fiducials in it to get to Tony Middleton. Expected yesterday
- Simulation of material in the interspace
- Action Item : John to give these files to the tracker group to make certain tracker resolution doesn't change
- Action Item : John to do a simulation with air in the interspace and He in the absorber volume
- Action Item : Pierrick/Mark - ensure that the pressure and material in the absorber volumes are recorded
- Action Item : Colin will check with Alan Bross whether the LiH absorber can withstand vacuum
- Cooling water discussion
- Action Item : Colin and John to talk and agree on the plan. Discussion complete but further work needed
- ISIS Monitoring of SS magnets
- Action Item : Josef/Colin to check with Alan about monitoring requirements for the SS magnets.
- BLOCS and SOCS
- Action Item :
Steve to remind Adam that we need SOCs.
- Action Item :
- DS PS Installation
- Action Item : Colin is to get an estimate for power supply installation time. 2 sets of 2-3 days over a space of 2 weeks
- Realignment work
- Upcoming and ongoing activities : chaired by DC
- Operations calendar
- C&M Network reconfiguration
- Realignment work schedule
- Tracker patch panel leak work
- Date of DSA closure and work required before hand
- Henry's January workplan
- Daresbury workplan
- AOB
- Date of next meeting : February 2nd
Minutes¶
Present: Andy, Vicky Bayliss, Paolo, Durga, Chris R, Steve B, Paul K, Trevor H, Jaroslaw, Mike C, Victoria, John N, Paul H, Henry N, Ken L
Review of action items¶
- Realignment work
- Action Item (Colin to prod John C) : need to get John's data with fiducials in it to get to Tony Middleton. Expected yesterday
- Done
- Action Item (Colin to prod John C) : need to get John's data with fiducials in it to get to Tony Middleton. Expected yesterday
- Simulation of material in the interspace
- Action Item : John to do a simulation with air in the interspace and He in the absorber volume
- John presented plots to show that there is little effect. However, these plots include systematic effects in the reconstruction
that are being worked on. - Action Item : John will look at the effect of air on the MC truth scattering angle
- John presented plots to show that there is little effect. However, these plots include systematic effects in the reconstruction
- Action Item : Pierrick/Mark - ensure that the pressure and material in the absorber volumes are recorded
- Action Item : DC to work with Pierrick to ensure this
- Action Item : Colin will check with Alan Bross whether the LiH absorber can withstand vacuum
- Alan says that the LiH absorber can withstand vacuum.
- Action Item : John to do a simulation with air in the interspace and He in the absorber volume
- Cooling water discussion
- Action Item : Colin and John to talk and agree on the plan.
- Discussion complete but further work needed
- Action Item : Colin and John to talk and agree on the plan.
- ISIS Monitoring of SS magnets
- Action Item : Josef/Colin to check with Alan about monitoring requirements for the SS magnets.
- Josef doesn't think we need to monitor to the level we are doing now. Proposed to stop the 4-hour monitoring and just leave
the buzzer in the control room. - Action Item : Steve to talk to Alan Stevens about this possibility
- Josef doesn't think we need to monitor to the level we are doing now. Proposed to stop the 4-hour monitoring and just leave
- Action Item : Josef/Colin to check with Alan about monitoring requirements for the SS magnets.
- BLOCS and SOCS
- Action Item : Steve to remind Adam that we need SOCs.
- Done
- Action Item : Steve to remind Adam that we need SOCs.
- DS PS Installation
- Action Item : Colin is to get an estimate for power supply installation time.
- 2 sets of 2-3 days over a space
- Action Item : Colin is to get an estimate for power supply installation time.
Updates and future schedule¶
- Trevor outlined the DS PS return schedule. The PS is expected to return in mid-February (8th - 15th). He plans to start commissioning
the new PS on Feb 22nd. This could roll into the next week.
Updated by Boyd, Steven about 7 years ago · 16 revisions